I need help with this problem.
Solve by the elimination method
4x-3y=39
3x+2y=-9
What is the solution of the system
Type an ordered pair
Simplify your answer

Answer by checkley77(12844)

4x-3y=39 MULTIPLY BY 2
3x+2y=-9 MULTIPLY BY 3 & ADD.
8X-6Y=78
9X+6Y=-27
-----------------------
17X=51
X=51/17
X=3 ANS.
4*3-3Y=39
12-3Y=39
-3Y=39-12
-3Y=27
Y=27/-3
Y=-9 ANS.
PROOF:
3*3+2*-9=-9
9-18=-9
-9=-9